OpenStudy (viixxen):

Verify the basic identity. What is the domain of validity? cot θ = cos θ csc θ I already did the basic identity part, so I just need help with finding the domain of validity. Thank you! :)

OpenStudy (jdoe0001):

\(\bf cot(\theta)=\cfrac{cos(\theta)}{sin(\theta)}\impliedby \textit{valid domain is when is not "undefined"}\) meaning, when the sine function for that angle isn't "0" so, that'd be the valid domain for it :)

jhonyy9 (jhonyy9):

cot theta = cos theta /sin theta csc theta = 1/sin theta cos theta*csc theta = cos theta *1/sin theta = cos theta / sin theta = cot theta cot theta = cot theta